OpenStudy (anonymous):

given f(x) = 3x-1, find f(x)-f(1) / (x-1)

OpenStudy (anonymous):

is the equation written like this? (3x-1)-1 / ( 3x-1 ) -1 ?

zepdrix (zepdrix):

I think you want to write it like this :) (3x-1)-1 / x - 1 The bottom x should only be the argument of the function :D

OpenStudy (anonymous):

Actually, it is written as \[\frac{ 3x - 1 - (3(1) - 1) }{ x - 1 }\]

OpenStudy (anonymous):

what do you mean by the only agrument in the function?

zepdrix (zepdrix):

Oh f(1) ah sorry about that D:

OpenStudy (anonymous):

Saying, f(x) = 3x -1, right? So f(1) = 3(1) - 1. Basically you replace the x to 1.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

oops, i overlooked the fact that the denom didn't have an f. thanks!

OpenStudy (anonymous):

No problem.